OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Show work please! Thanks! x - y = 3 and x + y = 7

OpenStudy (saifoo.khan):

C the chat.

OpenStudy (anonymous):

x-y=3 +x+y=7 --------- 2x=10 x=5 x-y=3 5-y=3 y=2